OpenStudy (anonymous):

how many solutions are there to 14x+2=12x

OpenStudy (cwrw238):

linear equation one solution can you solve it?

OpenStudy (anonymous):

14x + 2 = 12x -12x -12x 2x +2 = 0 -2 -2 2x = -2 do you know how to solve from here???

OpenStudy (anonymous):

if you don't then here is the rest: 2x/2 = -2/2 x = -1